# Geometrical frustration, power law tunneling and non-local gauge fields from scattered light

### v2

@@ Abstract
The versatility characteristic of our approach becomes even more spectacular when considering rotations of the hoppings in the complex plane. Here, as shown in Fig. 4 artificial gauge fields take place. Contrary to other setups like cold atomic [18, 14], solid state [23], photonic [26] and optomechanical [29] settings, our setup allows for arbitrary adjusting of the Peierls phase by one control parameter – the relative orientation between the molecular cloud and the LG modes. More specifically, the Peierls phase acquired by the photon hopping from the $n-$th LG mode to the $(n+k)-$th LG mode is given by the sum of the azimutal phase difference between the two modes and the phase $\varphi_{k}$. The ability to tune $\varphi_{k}$ arbitrarily by simply rotating the corresponding $\cos$-component in the density of the molecular cloud (see Eq. (Manipulation of couplings by geometrical shaping:—)) around the $z$-axis, allows us to engineer arbitrary translationally invariant Peierls phases for each hopping independently. In Fig. 4 (a) we show the complex phase of the kinetic part of the effective Hamiltonian, chosen so that in the triangular ladder geometry, a phase of $\Theta_{1}=\pi$ is acquired by going around a plaquette. In Fig. 4 (b) we add one layer of complexity by introducing a next-to-next nearest neighbor hopping with a complex amplitude, so that two different phases – $\Theta_{1}=\pi$ and $\Theta_{2}=\pi/2$ are acquired, dependent on the triangle. Specifically here the generated gauge fields can become effectively long-range as they can extended over different triangular plaquettes. As we underline, this unique feature can be made by controlling the density profile of the molecular cloud thus further proving the conceptual simplicity of our designed setup.
Conclusions and outlook.—
@@
We studied the behavior of photons scattering off-resonantly from a molecular cloud. In such a context, we derived a highly versatile Bose-Hubbard model where the elementary constituents are represented by the scattered photonic modes. Here, we demonstrated that the molecular distribution represents a powerful control parameter in order to achieve interesting configurations. Specifically, we first proved that, by properly shaping the molecular geometry, geometrical frustration naturally emerges. In this regard, we remark that configurations where geometrical frustration coexists with local and non-local interactions, as in our model, have been recently proposed as the ideal playground to explore states of matter characterized by chiral order [50] and exotic phase transitions [49]. Inspired by recent results proving that power law tunneling processes with small decay exponent might represent a fundamental tool for quantum computing purposes [33, 34], we further demonstrated that long-range hopping amplitudes and their relative decay can be shaped at will. Finally, we showed that for rotations of the molecular cloud both local and non-local gauge fields are acquired by scattered modes at each tunneling event. Crucially, this feature represents a fundamental tool to create paradigmatic topological phases [61].
@@
We conclude by stressing that our work naturally motivates promising research directions. In particular, the search for specific molecules with a small gap and well-separated ground and first excited vibrational state from the rest of the spectrum can in principle open the path towards the exploration of strongly interacting photonic regimes [62, 63]. Our results thus represent an alternative an powerful approach to unveil novel quantum regimes induced by complex coupling structures.
Acknowledgments.—
@@
We warmly thank S. Asban for his contribution in the initial phase of this project. We also thank I. Carusotto and O. Firstenberg for discussions. ICFO-QOT group acknowledges support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033,  CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR” (PRTR-C17.I1), FPI); QUANTERA MAQS PCI2019-111828-2;  QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168628), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project; European Union’s Horizon 2020 program under the Marie Sklodowska-Curie grant agreement No 847648; P.P.P. acknowledges also support from the “Secretaria d’Universitats i Recerca del Departament de Recerca i Universitats de la Generalitat de Catalunya” under grant 2024 FI-3 00390, as well as the European Social Fund Plus. L. B. acknowledges financial support within the DiQut Grant No. 2022523NA7 funded by European Union– Next Generation EU, PRIN 2022 program (D.D. 104- 02/02/2022 Ministero dell’Università e della Ricerca).
The data set used in this paper is available on Zenodo [64].

### v3

@@ Abstract
The versatility characteristic of our approach becomes even more spectacular when considering rotations of the hoppings in the complex plane. Here, as shown in Fig. 4 artificial gauge fields take place. Contrary to other setups like cold atomic [22, 4], solid state [64], photonic [31] and optomechanical [65] settings, our setup allows for arbitrary adjusting of the Peierls phase by one control parameter–the relative orientation between the molecular cloud and the LG modes. More specifically, the Peierls phase acquired by the photon hopping from the $n$th LG mode to the $(n+k)$th LG mode is given by the sum of the azimuthal phase difference between the two modes and the phase $\varphi_{k}$. The ability to tune $\varphi_{k}$ arbitrarily by simply rotating the corresponding $\cos$ component in the density of the molecular cloud [see Eq. (Manipulation of couplings by geometrical shaping—)] around the $z$ axis, allows us to engineer arbitrary translationally invariant Peierls phases for each hopping independently. In Fig. 4 (a) we show the complex phase of the kinetic part of the effective Hamiltonian, chosen so that in the triangular ladder geometry, a phase of $\Theta_{1}=\pi$ is acquired by going around a plaquette. In Fig. 4 (b) we add one layer of complexity by introducing a next-to-next-nearest neighbor hopping with a complex amplitude, so that two different phases – $\Theta_{1}=\pi$ and $\Theta_{2}=\pi/2$ are acquired, dependent on the triangle. Specifically here the generated gauge fields can become effectively long range as they can extended over different triangular plaquettes. As we underline, this unique feature can be made by controlling the density profile of the molecular cloud thus further proving the conceptual simplicity of our designed setup.
Conclusions and outlook—
@@
We studied the behavior of photons scattering off resonantly from a molecular cloud. In such a context, we derived a highly versatile Bose-Hubbard model where the elementary constituents are represented by the scattered photonic modes. Here, we demonstrated that the molecular distribution represents a powerful control parameter in order to achieve interesting configurations. Specifically, we first proved that, by properly shaping the molecular geometry, geometrical frustration naturally emerges. In this regard, we remark that configurations where geometrical frustration coexists with local and nonlocal interactions, as in our model, have been recently proposed as the ideal playground to explore states of matter characterized by chiral order [16] and exotic phase transitions [8]. Inspired by recent results proving that power law tunneling processes with small decay exponent might represent a fundamental tool for quantum computing purposes [44, 41], we further demonstrated that long-range hopping amplitudes and their relative decay can be shaped at will. Finally, we showed that for rotations of the molecular cloud both local and nonlocal gauge fields are acquired by scattered modes at each tunneling event. Crucially, this feature represents a fundamental tool to create paradigmatic topological phases [33].
@@
We conclude by stressing that our work naturally motivates promising research directions. In particular, the search for specific molecules with a small gap and well-separated ground and first excited vibrational state from the rest of the spectrum can in principle open the path toward the exploration of strongly interacting photonic regimes [21, 18]. Our results thus represent an alternative and powerful approach to unveil novel quantum regimes induced by complex coupling structures.
Acknowledgments—
@@
We warmly thank S. Asban for his contribution in the initial phase of this project. We also thank I. Carusotto and O. Firstenberg for discussions. ICFO-QOT group acknowledges support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033,  CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR” (PRTR-C17.I1), FPI); QUANTERA MAQS PCI2019-111828-2;  QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme cofunded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT U16-011424, cofunded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168628), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project; European Union’s Horizon 2020 program under the Marie Sklodowska-Curie grant agreement No 847648; P.P.P. acknowledges also support from the “Secretaria d’Universitats i Recerca del Departament de Recerca i Universitats de la Generalitat de Catalunya” under grant 2024 FI-3 00390, as well as the European Social Fund Plus. L. B. acknowledges financial support within the DiQut Grant No. 2022523NA7 funded by European Union– Next Generation EU, PRIN 2022 program (D.D. 104- 02/02/2022 Ministero dell’Università e della Ricerca).
Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them.
The data that support the findings of this article are openly available [51].